# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ 4.2.2 Different Domains of DAP
Architecture expansion has proven effective in acquiring robust coding abilities and preventing vertical forgetting simultaneously. IRCoder [221] investigates the potential of utilizing compiler intermediate representations to enhance the multilingual transferability of Code LLMs. By conducting DAP on code grounded in intermediate representations with LoRA [109], IRCoder achieves superior multilingual programming instruction following, enhanced multilingual code understanding, and increased robustness to prompt perturbations. Llama Pro [310], initialized from Llama2, undergoes continuous pre-training on a combination of code and math data. It expands the original Llama2 architecture by dynamically adding multiple identity copies of the transformer blocks. These added blocks initially serve as identity mappings, preserving the original functionality, and will be tuned to adapt to the DAP data. The proposed expansion method is asserted to be more resilient against catastrophic forgetting of previous general knowledge compared to sequential fine-tuning and other parameter-efficient tuning methods like LoRA.
@@
The three aforementioned works highlight the importance of domain-adaptive pre-training in the realm of code. However, it is crucial to note that the problem definition and conventional architectures of existing Code LLMs may present compatibility challenges for DAP deployment. This situation leads to trade-offs between specialization in the code domain and generalization ability in the natural language domain.
@@
Other Domains. ECONET [93] enhances the model’s ability to reason about event temporal relations through a dedicated continual pre-training phase. In this phase, temporal and event indicators are masked out, and a contrastive loss is applied to the recovered masked tokens. Results demonstrate that incorporating this domain-adaptive pre-training stage significantly improves performance on final tasks compared to direct fine-tuning. In Concept-Aware Language Model (CALM) [369], the authors introduce a data-efficient domain-adaptive pre-training approach. This method incorporates both generative and discriminative commonsense reasoning to enhance the concept-centric commonsense reasoning ability of large language models. Similar to ECONET, CALM’s DAP phase is specifically tailored for concept-centric reasoning tasks. Consequently, even a small number of data examples for continued pre-training can lead to notable improvements in downstream tasks.
@@
EcomGPT-CT [187] is a large language model tailored for the E-commerce through domain-adaptive pre-training. Given that E-commerce data often exhibits a semi-structured format, stored in tables or databases, EcomGPT-CT employs a data mixing strategy. This approach transforms semi-structured data into a set of nodes and edges, samples a cluster of nodes, and then extracts and concatenates them into a training example. During DAP, it combines the general-domain corpus with E-commerce data at a ratio of 2:1, which is significantly lower than the common setting adopted by other works.
Notably, AdaptLLM [49] challenges the data efficiency of conventional DAP by adopting a novel approach inspired by human learning patterns in reading comprehension. Following techniques from previous work [281], it transforms raw corpora into (raw text, question, answer) format, creating intrinsic tasks. The model is trained on reading tasks (next token prediction on the raw text) and comprehension tasks (question-answering based on the corpora). During DAP, to ensure instruction diversity beyond predefined templates, it integrates substantial amounts of general instruction tuning data at ratios ranging from 1:1 to 1:2 across various domains like biomedicine, finance, and law. Compared to traditional DAP approaches that utilize raw domain-specific corpora as-is, AdaptLLM’s method demonstrates superior domain-specific knowledge adaptation and minimal vertical forgetting.
4.3 Continual Fine-Tuning (CFT)
@@
Background of Continual Fine-Tuning (CFT). Continual Fine-Tuning (CFT) lies at the bottom layer of the vertical continuity, where models are trained on successive homogeneous tasks drawn from an evolving data distribution. As the service-oriented layer of LLM, it doesn’t require consideration of further adaptation to another downstream tasks, simplifying optimization objectives to a great extent: better adaptation and less forgetting. In essence, the goal is to improve the overall performance. The challenge of CFT has been extensively explored in the continual learning community. Essentially, all existing continual learning literature can be seen as a variant of CFT: models are either randomly initialized or initialized from pre-trained weights and undergo CFT thereafter. In the realm where continual learning intersects with natural language processing, we will only briefly outline the most notable works in this domain in 4.3.2, and direct interested readers to additional survey literature on this topic [22, 132].
In the era of LLMs, new computational paradigms in CFT have emerged and attracted significant attention within the research community. These topics include Continual Instruction Tuning (CIT, Section 4.3.3), Continual Model Refinement (CMR, Section 4.3.4), Continual Model Alignment (CMA, Section 4.3.5), and Continual Learning for Multimodal Language Models (CMLLMs, Section 4.3.6). While all these fall under the umbrella of CFT, each presents distinct features and challenges. In CIT, models must generalize to new tasks encoded in instructions, requiring semantic understanding [358]. CMR demands fine-grained, possibly example-level, operations for model refinement, differing from task-based approaches [96]. CMA aligns models with evolving human preferences, challenging due to subjective nature and lack of clear task boundaries [165, 352]. In CMLLMs, addressing the composite architectural design and preventing catastrophic forgetting are key challenges [100, 213]. Detailed exploration of these sub-categories follows in subsequent chapters.
@@
Summary of Continual Fine-Tuning LLMs. We have organized existing research on continual fine-tuning in Table 3, categorizing studies into sub-categories: (i) general continual fine-tuning, (ii) continual instruction tuning (CIT), (iii) continual model refinement (CMR), (iv) continual model alignment (CMA), and (v) continual multimodal LLMs (CMLLMs). The table includes details on incremental learning types (X-IL), LLM architecture, and employed continual learning techniques and evaluation metrics. After discussing general observations on CFT in Section 4.3.1, we’ll delve into each sub-category in detail.
@@
Table 3: Summary of the existing studies on Continual Fine-Tuning LLMs, where the papers are organized in five main categories based on what downstream tasks they are designed to tackle, including (i) General Continual Fine-Tuning (CFT); (ii) Continual Instruction Tuning (CIT); (iii) Continual Model Refinement (CMR); (iv) Continual Model Alignment (CMA); (v) Continual Multimodal LLMs (CMLLMs), which is shown in the column of CFT Type. The column of X-IL shows what continual learning paradigm the study includes [282], where TIL represents task-incremental learning, meaning task ID/information is provided during inference; DIL represents domain-incremental learning, meaning the tasks are defined in the same format, and no task ID/information is available during inference; CIL represents class-incremental learning, meaning the task ID needs to be further inferred when testing. Among 34 papers shown in the table, 100% (34/34) of them explicitly deploy the continual techniques to address the challenge of CFT. Furthermore, 30% (10/34) of them develop their own new techniques that cannot be easily categorized into the three mainstream sets of continual learning algorithms.
@@
General Continual Fine-Tuning CTR [133] DIL    CIL BERT Adapter
[274] TIL BERT S-Replay
CIRCLE [342] DIL T5 Replay EWC Prompt
ConPET [268] DIL Llama Replay LoRA
[12] DIL    CIL BERT G-Prompt
[183] TIL DistilBERT ALBERT RoBERTa ER    DER    LwF
SEQ [365] TIL    CIL BERT Pythia GPT2 P-Freeze Tricks for Classfiers
LFPT5 [230] DIL T5 P-Replay
[306] DIL RoBERTa GPT2 Replay EWC     SI     RWalk
LR ADJUST [307] DIL XLM-R LR Scheduling
C3 [47] TIL T5 KD Prompt Tuning
CT0 [255] TIL T0 S-Replay
RCL [292] TIL LLaMA Vicuna Baichuan Replay
DynaInst [205] TIL BART Replay
CITB [358] TIL T5 Replay  AGEM L2    EWC AdapterCL
SSR [113] TIL LLaMA Alpaca RandSel KMeansSel
@@
KPIG [103] DIL     TIL LLaMA Baichuan DynaInst PCLL  DCL L2    EWC DARE  LM-Cocktail KPIG
ConTinTin [337] TIL BART Replay InstructionSpeak
O-LoRA [291] TIL LLaMA Alpaca O-LoRA
Continual Instruction Tuning SAPT [363] TIL T5     LLaMA SAPT
InsCL [294] TIL LLaMA Replay InsCL
CMR [162] DIL BART ER    MIR    MLR L2    EWC
GRACE [96] DIL T5   BERT   GPT2 Adapter
WilKE [108] DIL GPT2 GPT-J Adaptor
Larimar [58] DIL BERT GPT-J Kanerva Memory
MELO [339] DIL BERT GPT2    T5 LoRA
Continual Model Refinement CME [156] DIL BERT Replay Inner-Prod. Reg.
COPF [351] TIL    DIL LLaMA Replay Function Reg. Prompt
AMA [165] DIL OpenLLaMA  Mistral Replay L1/L2 LoRA Adaptive Model Avg.
Continual Model Alignment CPPO [352] TIL GPT2 Weighting Strategy Prompt
EProj [100] TIL InstructBLIP TSIR Projector Exp.
Fwd-Prompt [364] TIL InstructBLIP BLIP2 Projector Exp.
CoIN [41] TIL LLaVA MoE  LoRA
Model Tailor [371] TIL InstructBLIP LLaVA-1.5 Model Tailor
Continually Fine-tuning Multimodal LLMs RebQ [362] TIL ViLT Prompt Tuning
4.3.1 General Observations about Continual Fine-Tuning
Examining the landscape of continual learning in the context of large language models (LLMs), and combined with the results shown in Table 3, we can discern significant trends and shifts within the research community. Firstly, there has been a noticeable transition in focus from class-incremental learning to domain-incremental and task-incremental learning paradigms. For example, in 12 papers of general continual fine-tuning (the first white section), only 3 papers study continually fine-tuning language models in the setting of class-incremental learning. In the remaining four topics, CIT completely belongs to the field of task-incremental learning, as the instruction provided to the models can be seen as a soft encoding of the task information; CMR completely belongs to the field of domain-incremental learning, as different editing examples follows the same problem definition, and no task information is provided during inference. In CMA and CMLLMs, no examples of class-incremental learning has been reported.
@@ 4.3.1 General Observations about Continual Fine-Tuning
It has been a longstanding common sense in the continual learning community that class-incremental learning, as it requires the model to predict the context label and within-context label at the same time [282, 288, 139], is the most challenging continual learning scenario and hence receives most of the attention from the community. However, the evolution observed so far suggests a recognition of the broader spectrum of challenges faced in real-world applications of continually learning LLMs. This growing awareness of the importance of task-incremental and domain-incremental learning underscores the necessity for more comprehensive and specific approaches to continual LLMs, with precedent continual learning literature provided as reference.
Furthermore, in continual fine-tuning (CFT), continual learning techniques enjoy broader adoption and explicit exploration compared to CPT and DAP. In Table 3, all 34 papers explicitly deploy the continual techniques, 50% of which develop their own new techniques that cannot be easily categorized into the three mainstream sets of continual learning algorithms. For example, in instruction tuning, SAPT designs a shared attentive learning framework to enable the catastrophic forgetting mitigation and knowledge transfer at the same time [363]; in model refinement, Larimar proposes to adopt an external memory system, Kanerva Memory, that supports operations of read, write, and generate [58]; in model alignment, AMA proposes adaptive model averaging method to achieve Pareto-optimal when averaging different models for a reward-tax trade-off (alignment-performance trade-off) [165]; in continual learning multimodal LLMs, [362] devises a novel method named Reconstruct before Query (RebQ), harnessing the multi-modal knowledge from a pre-trained model to reconstruct the absent information for the missing modality.
Hence we can further conclude that, beyond mere replication of existing techniques, researchers are actively developing tailored solutions for various scenarios where an LLM needs to be continually updated directly for downstream tasks. This emphasis underscores the recognition of continual learning as a pivotal component in the development of resilient and adaptive language models, and further signals a maturation of the field towards more specialized and effective continual learning methodologies for large language models.
4.3.2 General Continual Fine-Tuning (General CFT)
@@
Researchers have long investigated the phenomenon of forgetting resilience in pre-trained large language models when fine-tuned for various downstream tasks [133, 274, 183, 365, 195]. Although the pre-trained weights initially position the model in a flat-loss basin, aiding adaptation to future tasks without severely impacting previous ones [95, 208, 199, 195], zero or near-zero forgetting is only observed at the representation level. This implies that while the model retains its ability to distinguish between task-specific representations, it may still forget specific task details [313, 274, 183, 365]. Therefore, additional measures are necessary when deploying these models in real-world applications [133, 342, 12, 230, 306, 47]. For instance, [306] demonstrate that relying solely on the intrinsic anti-forgetting and generalization abilities of LLMs can be risky in scenarios with changing code and API distributions, and simple baseline continual learning techniques can significantly enhance performance.
@@
Many studies advance beyond naive sequential fine-tuning, leveraging the inherent anti-forgetting nature of LLMs while avoiding overly complex techniques of continual learning. They favor sequential fine-tuning due to its simplicity and predictability in real-world scenarios [307, 365]. For instance, LR ADJUST [307] proposes a straightforward yet effective method of dynamically adjusting the learning rate to mitigate the overwriting of knowledge from new languages onto old ones. Building on the innate anti-forgetting ability of large language models like Pythia [21], SEQ [365] introduces several strategies for fine-tuning LLMs on a sequence of downstream classification tasks: (i) freezing the LLMs and old classifiers after the warm-up phase and learning new tasks, respectively; (ii) employing (cosine) linear classifiers when old data is (not) available; and (iii) pre-allocating future classifiers.
@@
Given the minimal forgetting observed at the representation level in continual learning, some studies aim to tackle the misalignment between the representation space and the decision-making layers, such as the final classification layer, by introducing representation-level constraints during continual fine-tuning. NeiAttn [12] exemplifies this approach by formulating classification tasks as masked language modeling and proposing a neighboring attention mechanism to counteract negative representation drift. This method seamlessly complements existing continual learning techniques [38, 37, 270], as its focus is independent of mainstream continual learning methods.
@@
Another line of approaches delves into refining the input/output format and network architectures of pre-trained language models, crafting specific structures and training methods to tackle the challenges of continual fine-tuning. For instance, CTR [133] integrates a pre-trained model at the onset of incremental learning, incorporating two CL-plugin modules within transformer layers. These modules consist of a task-specific module (TSM) for acquiring task-specific knowledge and a knowledge-sharing module (KSM) for selectively transferring previously learned similar knowledge. CIRCLE [342] manually designs diverse prompt templates for various types of buggy code, unifying them into representations solvable through cloze tasks. It employs difficulty-based example replay to enhance continual program repair, outperforming existing automatic program repair (APR) methods. LFPT5 [230] addresses lifelong few-shot language learning by consolidating sequence labeling, text classification, and text generation tasks into a text-to-text format. The model, treated as a generative model, undergoes prompt tuning and data generation, generating pseudo-examples from previous domains during adaptation to new tasks and performing knowledge distillation. In [357], the authors propose a method for adaptively adding compositional adapters during continual sequence generation tasks. Before training on new domains, a decision stage determines which trained module can be reused using hidden state mixing. During training, this module addresses the task of fine-tuning while simultaneously regenerating examples. C3 [47] merges parameter-efficient fine-tuning (PEFT) and in-context learning (ICL) in a teacher-student framework. The teacher model undergoes in-context tuning focused solely on the current domain, while the student model, equipped with tunable prompts, minimizes the KL-divergence of both models’ autoregressive token distributions per step to retain the teacher model’s few-shot capability.
4.3.3 Continual Instruction Tuning (CIT)

### v2

@@ 4.2.2. Different Domains of DAP
As the only work investigated so far that utilizes the general data replay to mitigate vertical forgetting in the code domain, Code Llama (rozière2024code, ) introduces a sophisticated training framework tailored for various coding tasks and model sizes. Initialized from Llama 2 weights, these models undergo DAP on a dataset composed of deduplicated public code, discussions about code, and a subset of natural language data. This mix of natural language data serves as a form of pseudo-replay to maintain the models’ proficiency in understanding natural language. Besides replay, architecture expansion has proven effective in acquiring robust coding abilities and preventing vertical forgetting simultaneously. IRCoder (paul2024ircoder, ) utilizes compiler intermediate representations to enhance the multilingual transferability of Code LLMs. By conducting DAP on code grounded in intermediate representations with LoRA (hu2021lora, ), IRCoder achieves superior multilingual programming instruction following, enhanced multilingual code understanding, and increased robustness to prompt perturbations. Llama Pro (wu2024llama, ) undergoes DAP on a combination of code and math data. It expands the original Llama2 architecture by dynamically adding multiple identity copies of the transformer blocks. These added blocks initially preserves the original functionality, and will be tuned for DAP. The proposed expansion method is shown to be more resilient against vertical forgetting compared to other parameter-efficient tuning methods like LoRA.
@@
The three aforementioned works highlight the importance of DAP for code LLMs. However, it is crucial to note that the problem definition and conventional architectures of existing Code LLMs may present challenges of compatibility for DAP deployment, and need to be addressed in the future.
@@
Other Domains. ECONET (han2021econet, ) enhances the model’s ability to reason about event temporal relations through a dedicated DAP phase. Temporal and event indicators are masked out, and a contrastive loss is applied to the recovered masked tokens. Results demonstrate that incorporating this DAP stage significantly improves performance on final tasks compared to direct fine-tuning. Concept-Aware Language Model (CALM) (zhou2020pre, ) introduces a data-efficient DAP approach for enhancing the concept-centric commonsense reasoning ability of LLMs. It incorporates both generative and discriminative commonsense reasoning tasks specifically tailored for concept-centric reasoning tasks. Consequently, even a small number of data examples for DAP can lead to notable improvements for downstream tasks.
@@
Aurora-M (nakamura2024aurora, ) and Swallow (fujii2024continual, ) adopt the simple replay strategy that mixes in a small portion of general data during DAP for their multi-lingual ability. Furthermore, Sailor (dou2024sailor, ) studies the optimal strategy of data mixing for DAP, balancing the general knowledge and capacity of different languages. EcomGPT-CT (ma2023ecomgptct, ) employs a data mixing strategy for DAP which transforms semi-structured E-commerce data into a set of nodes and edges, samples a cluster of nodes, and then extracts and concatenates them into a training example. It combines the general-domain corpus with E-commerce data at a ratio of 2:1, which is significantly lower than the common setting adopted by other works.
Notably, there are some papers studying other effective ways of DAP. AdaptLLM (cheng2024adapting, ) transforms raw corpora into (raw text, question, answer) format, creating intrinsic reading comprehension tasks. AdaptLLM demonstrates superior domain-specific knowledge adaptation and minimal vertical forgetting, thereby challenging the data efficiency of conventional DAP. Tag-LLM (shen2024tag, ) re-purposes the general-domain LLM into domain-specific one by multi-stage training of domain tags and function tags, without modifying the base LLM’s weights and thereby mitigates forgetting.
4.3. Continual Fine-Tuning (CFT)
@@
Background of Continual Fine-Tuning (CFT). Continual Fine-Tuning (CFT) lies at the bottom layer of the vertical continuity, where models are trained on successive homogeneous tasks drawn from an evolving data distribution. As the service-oriented layer of LLM, it doesn’t require consideration of further adaptation to another downstream tasks, simplifying optimization objectives to a great extent: better adaptation and less forgetting [^3] We direct interested readers to additional survey literature on the topic of general CFT (biesialska2020continual, ; ke2023continual, ).. In the era of LLMs, new computational paradigms in CFT have emerged and attracted significant attention within the research community. These topics include:
  • Continual Instruction Tuning (CIT), where models must generalize to new tasks encoded in instructions, requiring semantic understanding (zhang2023citb, ) (Section 4.3.3).
  • Continual Model Refinement (CMR), where fine-grained, possibly example-level solutions are required, differing from task-level approaches (hartvigsen2023aging, ) (Section 4.3.4).
  • Continual Model Alignment (CMA), which aligns models with evolving human preferences, challenging due to subjective nature and lack of clear task boundaries (lin2024mitigating, ; zhangcppo, ) (Section 4.3.5).
  • Continual Learning for Multimodal Language Models (CMLLMs), where addressing the composite architectural design and preventing catastrophic forgetting are key challenges (he2023continual, ; ni2023continual, ) (Section 4.3.6).
@@
We summarize existing studies on CFT in Table 3, categorizing studies into sub-categories as listed above. The table includes details on incremental learning types (X-IL), LLM architecture, and employed CL techniques and evaluation metrics. After discussing general observations on CFT in Section 4.3.1, we’ll delve into each sub-category in detail.
@@
Table 3. Summary of the existing studies on Continual Fine-Tuning LLMs, where the papers are organized in five main categories based on what downstream tasks they are designed to tackle, including (i) General Continual Fine-Tuning (CFT); (ii) Continual Instruction Tuning (CIT); (iii) Continual Model Refinement (CMR); (iv) Continual Model Alignment (CMA); (v) Continual Multimodal LLMs (CMLLMs), which is shown in the column of CFT Type. The column of X-IL shows what continual learning paradigm the study includes (van2022three, ), where TIL represents task-incremental learning, meaning task ID/information is provided during inference; DIL represents domain-incremental learning, meaning the tasks are defined in the same format, and no task ID/information is available during inference; CIL represents class-incremental learning, meaning the task ID needs to be further inferred when testing.
@@
General CTR (ke2021achieve, ) DIL | CIL BERT Adapter
(tao2022can, ) TIL BERT S-Replay
CIRCLE (wei2022circle, ) DIL T5 Replay EWC Prompt
ConPET (song2023conpet, ) DIL Llama Replay LoRA
(bai2023enhancing, ) DIL | CIL BERT G-Prompt
(luo2023investigating, ) TIL DistilBERT ALBERT | RoBERTa ER | DER | LwF
SEQ (zheng2023learn, ) TIL | CIL Pythia | BERT | GPT2 P-Freeze Tricks for Classifiers
LFPT5 (qin2021lfpt5, ) DIL T5 P-Replay
(weyssow2023usage, ) DIL RoBERTa | GPT2 Replay EWC | SI | RWalk
LR ADJUST (winata2023overcoming, ) DIL XLM-R LR Scheduling
C3 (chen2024parameterizing, ) TIL T5 KD Prompt Tuning
CT0 (scialom2022fine, ) TIL T0 S-Replay
RCL (wang2023trace, ) TIL LLaMA Vicuna | Baichuan Replay
DynaInst (mok2023large, ) TIL BART Replay
CITB (zhang2023citb, ) TIL T5 Replay | AGEM L2 | EWC AdapterCL
SSR (huang2024mitigating, ) TIL LLaMA | Alpaca RandSel | KMeansSel
@@
KPIG (he2024dont, ) DIL | TIL LLaMA | Baichuan DynaInst | PCLL | DCL L2 EWC DARE LM-Cocktail KPIG
ConTinTin (yin2022contintin, ) TIL BART Replay InstructionSpeak
O-LoRA (wang2023orthogonal, ) TIL LLaMA | Alpaca O-LoRA
CIT SAPT (zhao2024sapt, ) TIL T5 | LLaMA SAPT
InsCL (wang2024inscl, ) TIL LLaMA Replay InsCL
CMR (lin2022continual, ) DIL BART ER | MIR | MLR L2 | EWC
GRACE (hartvigsen2023aging, ) DIL T5 | BERT | GPT2 Adapter
WilKE (hu2024wilke, ) DIL GPT2 | GPT-J Adaptor
Larimar (das2024larimar, ) DIL BERT | GPT-J Kanerva Memory
MELO (yu2023melo, ) DIL BERT | GPT2 | T5 LoRA
CME (li2023continual, ) DIL BERT Replay Inner-Prod. Reg.
CMR WISE (wang2024wise, ) DIL GPT-J | Llama2 | Mistral Side Memory
COPF (zhang2023copf, ) TIL | DIL Llama Replay Function Reg. Prompt
AMA (lin2024mitigating, ) DIL OpenLLaMA | Mistral Replay L1 | L2 LoRA Adaptive Model Avg.
CMA CPPO (zhangcppo, ) TIL GPT2 Weighting Prompt
EProj (he2023continual, ) TIL InstructBLIP TSIR Projector Exp.
Fwd-Prompt (zheng2024antiforgetting, ) TIL InstructBLIP | BLIP2 Projector Exp.
CoIN (chen2024coin, ) TIL LLaVA MoE | LoRA
Model Tailor (zhu2024model, ) TIL InstructBLIP | LLaVA Model Tailor
CMLLMs RebQ (zhao2024reconstruct, ) TIL ViLT Prompt Tuning
4.3.1. General Observations on CFT
Examining the landscape of continual learning in the context of LLMs, and combined with the results shown in Table 3, we make several key observations about CFT.
@@ 4.3.1. General Observations on CFT
  • OBS-1: There has been a noticeable transition in focus from CIL to TIL and DIL. It has been a longstanding common sense in the CL community that CIL, as it requires the model to predict the context label and within-context label at the same time (van2022three, ; wang2024comprehensive, ; kim2022theoretical, ), is the most challenging CL scenario and hence receives most of the attention from the community. However, among all 35 papers presented in Table 3, only 3 papers study CFT of CIL. The transition of the research focus demonstrates the importance of TIL and DIL in the real-world applications of continual LLMs. More detail discussion of this transition will be included in Section 6.2.
  • OBS-2: In CFT, CL techniques enjoy broader adoption and explicit exploration compared to CPT and DAP. In Table 3, all 35 papers explicitly deploy the CL techniques, 50% of which develop new techniques that cannot be easily interpreted as trivial combination of existing classic CL techniques, e.g., shared attentive learning framework in SAPT (zhao2024sapt, ), external memory deployed in Larimar (das2024larimar, ), and adaptive model averaging method to achieve Pareto-optimal in AMA (lin2024mitigating, ), etc. This underscores the recognition of continual learning as a pivotal component in the development of resilient and adaptive LLMs.
4.3.2. General Continual Fine-Tuning (General CFT)
@@
Researchers have long investigated the phenomenon of forgetting resilience in pre-trained LLMs when fine-tuned for downstream tasks (ke2021achieve, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ), despite some discover the opposite (luo2023investigating, ). Although the pre-trained weights initially position the model in a flat-loss basin, aiding adaptation to future tasks without severely impacting previous ones (hao2019visualizing, ; neyshabur2020being, ; mirzadeh2022wide, ; mehta2023empirical, ), zero or near-zero forgetting is only observed at the representation level. This implies that while the model retains its ability to distinguish between task-specific representations, it may still forget specific task details (wu2021pretrained, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ). Therefore, additional measures are necessary when deploying these models in real-world applications (ke2021achieve, ; wei2022circle, ; bai2023enhancing, ; qin2021lfpt5, ; weyssow2023usage, ; chen2024parameterizing, ).
@@
Many studies advance beyond naive sequential fine-tuning, leveraging the inherent anti-forgetting nature of LLMs while avoiding the adoption of overly complex CL techniques (winata2023overcoming, ; zheng2023learn, ). For instance, LR ADJUST (winata2023overcoming, ) proposes a straightforward yet effective method of dynamically adjusting the learning rate to mitigate the overwriting of knowledge from new languages onto old ones. Building on the innate anti-forgetting ability of large language models like Pythia (biderman2023pythia, ), SEQ (zheng2023learn, ) introduces several strategies for fine-tuning LLMs on a sequence of downstream classification tasks, such as freezing the LLM and old classifier’s parameters after warm-up, and pre-allocating future classifiers, etc.
@@
Given the minimal forgetting observed at the representation level in CL, some studies aim to tackle the misalignment between the representation space and the decision-making layers by introducing representation-level constraints during CFT. NeiAttn (bai2023enhancing, ) exemplifies this approach by formulating classification tasks as masked language modeling and proposing a neighboring attention mechanism to counteract negative representation drift.
@@
Another line of approaches refines the input/output format and network architectures of pre-trained LLMs to be better suited for CFT. For instance, CTR (ke2021achieve, ) incorporates two CL-plugin modules that consist of a task-specific module (TSM) for acquiring task-specific knowledge and a knowledge-sharing module (KSM) for selectively transferring previously learned similar knowledge. CIRCLE (wei2022circle, ) manually designs diverse prompt templates for various types of buggy code, unifying them as the cloze task and employs difficulty-based replay to enhance continual program repair. LFPT5 (qin2021lfpt5, ) addresses lifelong few-shot language learning by consolidating sequence labeling, text classification, and text generation into a text-to-text generation task. It undergoes prompt tuning on generated pseudo-examples from previous domains when adapting to new tasks. In (zhang2022continual, ), the authors propose a method for adaptively adding compositional adapters during continual sequence generation tasks. Before training on new domains, a decision stage determines which trained module can be reused. During training, this module also regenerates examples of the past for replay. C3 (chen2024parameterizing, ) merges PEFT and in-context learning (ICL) in a teacher-student framework. The teacher model undergoes in-context tuning focused solely on the current domain, while the student model, together with tunable prompts, minimizes the KL-divergence between the output distribution and the ground truth and teacher model simultaneously.
4.3.3. Continual Instruction Tuning (CIT)